# VerifyBench: Benchmarking Reference-based Reward Systems for Large Language Models

### v2

@@ Appendix D Experimental Details
Appendix E LLM Usage
We list all the LLMs we used to generate completions for curated question in Table 6.
Appendix F Examples of VerifyBench
We provide some examples of VerifyBench with four different answer types in Figure 7, 8, 9 and 10.
Domain Source License # of Questions
general_reasoning BBH (Suzgun et al., 2022) MIT 4520
BBEH (Kazemi et al., 2025) Apache 2.0 6511
MMLU_pro (Wang et al., 2024) Apache 2.0 2000
natural_reasoning (Yuan et al., 2025) CC-BY-NC 4.0 10000
logic_reasoning AbductionRules (Young et al., 2022) MIT 1000
anlg (Bhagavatula et al., 2020) / 1000
anli (Nie et al., 2020) CC-BY-NC 4.0 1000
ARLSAT (Zhong et al., 2021) MIT 230
bAbI15 (Weston et al., 2015) / 1000
bAbI16 (Weston et al., 2015) / 1000
BoardgameQA (Kazemi et al., 2023) CC-BY-4.0 1000
clutrr (Sinha et al., 2019) CC-BY-NC 4.0 1000
FOLIO (Han et al., 2024) CC-BY-SA-4.0 134
hellaswag (Zellers et al., 2019) MIT 1000
logicbenchBQA (Parmar et al., 2024) MIT 1000
logicbenchMCQA (Parmar et al., 2024) MIT 1000
LogiQA (Liu et al., 2020) / 1000
MultiLogiEval (Patel et al., 2024) MIT 1000
NeuLRabductive (Xu et al., 2025a) / 1000
NeuLRdeductive (Xu et al., 2025a) / 1000
NeuLRinductive (Xu et al., 2025a) / 1000
ProntoQA (Saparov and He, 2023) Apache 2.0 500
ProofWriter (Tafjord et al., 2021) / 1000
ReClor (Yu et al., 2020) / 500
tablebench (Wu et al., 2025) Apache 2.0 886
math_reasoning AIME24 MIT 30
AIME25 MIT 30
asdiv-a (Miao et al., 2021) CC-BY-NC 4.0 1218
Math Odyssey (Fang et al., 2024) MIT 389
GPQA_diamond (Rein et al., 2023) MIT 198
gsm8k (Cobbe et al., 2021) MIT 1319
math401 (Yuan et al., 2023) / 392
mathematics (Saxton et al., 2019) Apache 2.0 3360
MATH(Hendrycks et al., 2021) MIT 5000
OlympiadBench-EN (He et al., 2024) MIT 675
SVAMP (Patel et al., 2021) MIT 1000
NuminaMath-CoT (Li et al., 2024) Apache 2.0 20000
orca-math-word-problems (Mitra et al., 2024) MIT 10000
ArtOfProblemSolving self-curated 7997
stackmathqa (Zhang, Yifan, 2024) CC-BY-4.0 10000
DeepMath-103K-RL (He et al., 2025) MIT 20000
Figure 4: Prompt for answer type classification.
Figure 5: Prompt for LLM-as-a-judge evaluation.
Figure 6: Prompt for LLM-as-a-judge evaluation without reference answers.

### v3

@@ H.2.1 Setup
H.2.2 Experimental Results
In our experiments, we applied rejection sampling to implement reference-based reward systems. For each question in the GSM8K and MATH training sets, we generated 64 candidate completions using Qwen2.5-Math-7B-Instruct (Yang et al., 2024) with a sampling temperature of 0.7. These responses were subsequently filtered by three verifier models with varying performance levels on VerifyBench: Llama-3.1-8B-Instruct, Qwen3-4B, and Qwen3-1.7B. Only completions consistently verified as correct were retained to form the SFT training data.
Appendix I Examples of VerifyBench
We provide some examples of VerifyBench with four different answer types in Figure 11, 12, 13 and 14. Due to space limitations, we omitted part of the text and used <omitted> as a placeholder.
Figure 11 presents a sample from the Numeric Values category, Figure 12 shows a sample from the Expression category, Figure 13 illustrates a sample from the Multi-choice category, and Figure 14 displays a sample from the String category.